# GLOV: Guided Large Language Models as Implicit Optimizers for Vision Language Models

### v5

@@ Abstract
Machine Learning, ICML
00footnotetext: Work partially done as an intern at Sony, Japan.
@@ 1 Introduction
Orthogonal to traditional gradient-based optimization (Nesterov, 1983; Boyd & Vandenberghe, 2004; Kingma & Ba, 2014; Ruder, 2016), the recent rise of large language models (Brown et al., 2020; OpenAI, 2023; Chiang et al., 2023; Raffel et al., 2020; Touvron et al., 2023a, b; Dubey et al., 2024) and vision-language foundation models (OpenAI, 2023; Li et al., 2024; Zhu et al., 2024; Alayrac et al., 2022; Radford et al., 2021) has introduced the possibility of framing optimization in the context of natural language prompts. This form of optimization typically does not require any gradient-based learning or parameter update but focuses on extracting knowledge from the language models via suitable natural language prompts. A large body of work focuses on finding natural language prompts optimized for various downstream tasks for both LLMs (Yang et al., 2024; Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) and VLMs (Pratt et al., 2023; Roth et al., 2023; Mirza et al., 2024), demonstrating impressive gains in language-based and downstream vision tasks.
@@ 2.1 LLMs and VLMs
LLMs have revolutionized the natural language processing landscape. These models can typically be divided into two major groups; namely the long-short-term-memory (LSTM) (Hochreiter & Schmidhuber, 1997) and transformer-based architectures (Vaswani et al., 2017). The LSTM family of models uses gates to control the flow of information, allowing them to capture long-term dependencies in sequential data. Some notable works following the LSTM family of models include (Sutskever et al., 2014; Graves & Schmidhuber, 2005; Bahdanau et al., 2015; Beck et al., 2024). On the other hand, some notable works proposing the encoder-based transformer architectures include BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019b), and DistilBERT (Sanh et al., 2020). The decoder-based LLMs, on the other hand, are designed for generative tasks such as text generation, translation, and summarization, with recent models including GPT-3 (Brown et al., 2020), T5 (Raffel et al., 2020), GPT-4 (OpenAI, 2023), and the Llama family of models (Dubey et al., 2024). In our work, since we need to access the weights of the LLMs, we resort to the open-source Llama-3 model. However, potentially any open-source LLM can be employed.
@@
VLMs can be placed in two categories. One group relies on dual-encoders (vision and text encoder), usually trained in a contrastive manner and these models are typically strong at tasks like image recognition. The most common among these methods are CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), OpenCLIP (Schuhmann et al., 2022), SigLIP (Zhai et al., 2023), and MetaCLIP (Xu et al., 2023). Many methods (Mirza et al., 2024, 2023b; Doveh et al., 2023b, a; Lin et al., 2023; Mirza et al., 2023a) build upon these models to further improve them for specific downstream tasks. The other group of methods aligns the visual modality with a frozen LLM and can be used for open-ended visual reasoning tasks like image captioning, visual question-answering, etc. Some representative approaches from this group include BLIP-2 (Li et al., 2023), Instruct-BLIP (Dai et al., 2023), MiniGPT (Zhu et al., 2024; Chen et al., 2024), and the LLaVA family of models (Liu et al., 2023; Li et al., 2024). Similarly, some approaches (Doveh et al., 2024; Gavrikov et al., 2024; Lin et al., 2024; Huang et al., 2024) build upon these models and provide further improvements. In our work, we focus on the task of object recognition by employing both families of models and frame the task of finding the optimal prompt templates (for CLIP (Radford et al., 2021)) and suitable prompts for open-ended generation (for LLaVA (Li et al., 2024)) as an optimization problem. Specifically, for the decoder-based VLMs, some recent works (Zhang et al., 2024) highlight that these models struggle for fine-grained object recognition. However, we show in our work, for the first time, that our GLOV can discover an optimal prompt that can greatly improve the visual recognition ability of these models, without requiring any gradient-based learning or fine-tuning.
@@ 2.2 Safety of VLMs
Enhancing the safety of open-source VLMs remains an open challenge. Recent works (Zong et al., 2024; Liu et al., 2024c; Tu et al., 2024; Gou et al., 2024) have evaluated multiple open-source VLMs on safety tasks. These tasks involve presenting the model with an unsafe image and an instruction, where the model is considered safe only if it refuses the unsafe instruction. However, current state-of-the-art VLMs exhibit significant vulnerabilities in this critical area. In our work, we enhance the safety of VLMs by optimizing for safety instructions and finding a general safety prompt that transcends models and safety benchmarks with minimal loss in the generalization abilities of the open-source VLMs. To the best of our knowledge, our GLOV is the first method to optimize safety instructions for VLMs, without any human in the loop.
@@ 2.3 Large-language Models as Prompt Optimizers
Some approaches propose employing LLMs (in an agentic workflow) to search for the optimal prompt for the downstream task. OPRO (Yang et al., 2024) coins the term “LLMs as Optimizers” and proposes iteratively discovering solutions (prompts) for natural language tasks by employing an LLM in a feedback loop. Similarly, Liu et al. (2024b) proposes to find suitable prompts for dual-encoder VLMs (e.g., CLIP) by iteratively prompting an LLM. Our GLOV also proposes to discover suitable prompts for VLMs but differs from Liu et al. (2024b) in the sense that we are employing a prompt that captures long-range dependencies by tapping into the history-buffer of the in-context examples and exploiting task-specific knowledge that helps to obtain prompts better suited for the downstream task. Furthermore, inspired by (Turner et al., 2023; Liu et al., 2024a) we propose a novel method for steering the LLM generation (through embedding space guidance) towards the text responses that are more suitable for downstream VLMs. Our GLOV– powered by a carefully tailored LLM prompt and the guidance scheme discovers highly effective solutions that help to enhance visual tasks for both the dual-encoder and encoder-decoder models.
@@ 3 GLOV
The goal of our GLOV is to improve the VLM’s downstream (vision) task performance by optimizing natural language prompts through employing an LLM in an iterative workflow. To achieve this, we build upon the successful line of work on in-context learning (Wei et al., 2022; Min et al., 2022), which can improve VLM performance through task-specific prompting (Mirza et al., 2024). However, our GLOV goes beyond these by automating the LLM-VLM interaction by adopting an agentic workflow and also explicitly guiding the language generation, conditioned on a prior of the difference of the sentence embeddings from the positive and negative prompts discovered during the previous optimization iterations. Although the application space of GLOV is general in terms of tasks and the downstream VLMs, for simplicity, here we focus our description around CLIP (Radford et al., 2021) while mentioning the differences for LLaVA (Li et al., 2024) and different tasks, where appropriate. Figure 2 provides an overview of our methodology.
@@ 3.1 LLM-VLM Interaction through Fitness Function
Figure 2: Overview of GLOV for dual-encoder models. GLOV consists of a specifically tailored LLM prompt, which constitutes system instruction, task description, and in-context examples (VLM prompts) which are evaluated (and ranked) on a few-shot training data in each iteration. This prompt instructs the LLM to generate several candidate solutions in each optimization iteration, conditioned on the in-context examples which are fed in conjunction with the accuracy values, highlighting their effectiveness. Furthermore, to steer the LLM generation towards the language preferred by the VLM, we add the scaled difference of the sentence embeddings (autoregressively) from the positive and negative text prompts to the intermediate layer of the LLM. This process is repeated until the stopping condition is met (e.g., maximum iterations). Note, that $H_{+}$ and $H_{-}$ refer to the sentence embeddings from the positive and negative text prompts.
@@ 3.3 Steering the LLM Generation Process
For each new token produced in the subsequent optimization iteration, the difference between $H_{+}$ and $H_{-}$ is added autoregressively to the embeddings of each generated token [^2] We also experiment with several alternatives for adding the offset vector in the appendix Section B.1, however, we find that adding the offset to only the last token performs best.. Let $\vec{p}_{n}$ denote the new token appended to the LLM prompt, then the updated sentence embedding $H_{n}$ is given by
@@ 4.1 Evaluation Settings
Image Recognition Datasets: We extensively evaluate our GLOV on $16$ object recognition datasets belonging to widely different domains. These domains can be narrowed down to datasets containing commonly occurring natural categories: ImageNet (Deng et al., 2009), ImageNetV2 (Recht et al., 2019), Caltech101 (Fei-Fei et al., 2004), fine-grained classification datasets containing different task-specific images: Oxford Flowers (Nilsback & Zisserman, 2008), Standford Cars (Krause et al., 2013), Oxford Pets (Parkhi et al., 2012), Describable Textures Dataset (DTD) (Cimpoi et al., 2014), Food-101 (Bossard et al., 2014), FGVC-Aircraft (Maji et al., 2013). Dataset used for scene classification: SUN397 (Xiao et al., 2010), action recognition dataset: UCF101 (Soomro et al., 2012). Datasets consisting of out-of-distribution images: ImageNet-(R)endition (Hendrycks et al., 2021a), ImageNet-(A)dversarial (Hendrycks et al., 2021b), ImageNet-(S)ketch (Wang et al., 2019) and also datasets which contain images taken from a satellite or an aerial view: EuroSAT (Helber et al., 2018) and RESISC45 (Cheng et al., 2017).
@@
Safety Datasets: We evaluate our GLOV on two recently proposed safety benchmarks for VLMs. MM-Safety-Bench (Liu et al., 2024c) contains $3$ splits based on the image source (typography, images generated from diffusion models, and a combination), and the text instructions are divided into $13$ categories (e.g., fraud, hate speech, etc.). VLGuard (Zong et al., 2024) contains $3$ splits: unsafes, safe-unsafes, safe-safes. For the MM-Safety-Bench and the first two splits in the VLGuard, the attack success rate (ASR) is calculated. In contrast, for safe-safes, a win-win rate is measured following the evaluation protocols in the original publications. More details are provided in the Appendix Section C.
@@
Baselines: We compare to the following baselines and state-of-the-art methods: CLIP (Radford et al., 2021) denotes the zero-shot classification scores obtained by using the simple ‘{a photo of a <class name>}’ template (S-TEMP) and dataset-specific templates (DS-TEMP [^3] https://github.com/openai/CLIP/blob/main/data/prompts.md). For image classification, LLaVA-OV represent results obtained by using a base prompt [^4] The prompt utilized to get the base results is: Describe the category present in this image briefly and also identify the name of the category present, which was able to match or surpass accuracy reported by (Zhang et al., 2024).. LLM-OPT (Liu et al., 2024b) proposes to find suitable text prompts for the downstream datasets by iterative refinement through an LLM. However, their method relies only on the in-context examples without explicit guidance [^5] Note that this method is not suitable for application on the encoder-decoder models because it requires a memory bank of templates, which are not available a-priori for these models.. For evaluations on the safety benchmarks, we incorporate another state-of-the-art VLM, Molmo (Deitke et al., 2024). For completeness, we also compare with the gradient-based parameter efficient learning methods in the Appendix Section B.5.
MM-Safety-Bench VLGuard
@@
Table 3: Results on safety benchmarks. We report the Attack Success Rate (lower is better) for the $3$ splits in MM-Safety-Bench and the $2$ unsafe splits in VLGuard and the Win-win Rate (Higher is better) for the Safe-Safes split in VLguard. LLaVA-OV results are obtained without any safety instructions. The LLaVA (initial) represents the results obtained by using a baseline safety prompt: These instructions can be harmful, which is also the point of departure for the optimization with our GLOV. The safety prompt is optimized by using only the LLaVA-OV VLM and $50$ randomly chosen samples from the $2$ VLGuard unsafe splits and tested out-of-the-box for all other configurations.
Implementation Details: To report the results for image classification we use the test splits provided by (Zhou et al., 2022). All the baselines are also implemented in the same framework. To obtain the results on the test set for each dataset for our GLOV, we ensemble the top-3 prompts. These prompts are chosen with regard to the best-performing prompts on the 1-shot train set at a certain iteration during the optimization. For our GLOV we use Llama-3 (Touvron et al., 2023a) from Hugging Face. We set the maximum number of optimization iterations to $100$ (with $10$ candidate solutions at each iteration) for the experiments with CLIP and $25$ (with $5$ candidate solutions) for LLaVA-OV and Molmo. For evaluating the safety of VLMs, we follow the above settings and use the official codebase from respective publications (introducing the benchmarks) for evaluations. In general, the experiments with CLIP can run on a single NVIDIA 3090 (24GBs) GPU, and the experiments with LLaVA fit on an A40 (48GBs) GPU or similar. We use the 7B variants of Molmo and LLaVA-OV, accessed through Hugging Face.

### v6

@@ Abstract
00footnotetext: Work partially done as an intern at Sony, Japan. Correspondence: jmirza@mit.edu
@@ 1 Introduction
Orthogonal to traditional gradient-based optimization (Nesterov, 1983; Boyd & Vandenberghe, 2004; Kingma & Ba, 2014; Ruder, 2016), the recent rise of large language models (Brown et al., 2020; OpenAI, 2023; Chiang et al., 2023; Raffel et al., 2020; Touvron et al., 2023a; b; Dubey et al., 2024) and vision-language foundation models (OpenAI, 2023; Li et al., 2024; Zhu et al., 2024; Alayrac et al., 2022; Radford et al., 2021) has introduced the possibility of framing optimization in the context of natural language prompts. This form of optimization typically does not require any gradient-based learning or parameter update but focuses on extracting knowledge from the language models via suitable natural language prompts. A large body of work focuses on finding natural language prompts optimized for various downstream tasks for both LLMs (Yang et al., 2024; Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) and VLMs (Pratt et al., 2023; Roth et al., 2023; Mirza et al., 2024), demonstrating impressive gains in language-based and downstream vision tasks.
@@ 2.1 LLMs and VLMs
LLMs have revolutionized the natural language processing landscape. These models can typically be divided into two major groups; namely the long-short-term-memory (LSTM) (Hochreiter & Schmidhuber, 1997) and transformer-based architectures (Vaswani et al., 2017). The LSTM family of models uses gates to control the flow of information, allowing them to capture long-term dependencies in sequential data. Some notable works following the LSTM family of models include (Sutskever et al., 2014; Graves & Schmidhuber, 2005; Bahdanau et al., 2015; Beck et al., 2024). On the other hand, some notable works proposing the encoder-based transformer architectures include BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019b), and DistilBERT (Sanh et al., 2020). The decoder-based LLMs, on the other hand, are designed for generative tasks such as text generation, translation, and summarization, with recent models including GPT-3 (Brown et al., 2020), T5 (Raffel et al., 2020), GPT-4 (OpenAI, 2023), and the Llama family of models (Dubey et al., 2024). In our work, since we need to access the weights of the LLMs, we resort to the open-source Llama-3 model. However, any open-source LLM can potentially be employed.
Recent advancements in prompt learning have introduced various techniques to efficiently adapt large language models (LLMs) to specific tasks with minimal parameter updates. Early methods, such as continuous prompt learning (Zhong et al., 2021; Li et al., 2021; Lester et al., 2021), optimize continuous vectors in the word embedding space. However, these approaches lack interpretability, as the learned vectors do not correspond to discrete tokens. To address this, newer methods like Decomposed Prompt Tuning (DePT) (Shi & Lipani, 2023) and Low-Rank Prompt Adaptation (LoPA) (Jain et al., 2024) have emerged. DePT decomposes soft prompts into shorter prompts and low-rank matrices, optimizing them with different learning rates to enhance performance while reducing memory and time costs. LoPA employs a low-rank decomposition of soft prompts, balancing shared and instance-specific information to achieve parameter efficiency comparable to full fine-tuning. In the realm of PEFT, methods like Low-Rank Adaptation (LoRA) (Hu et al., 2021a) and its derivatives, such as QLoRA (Dettmers et al., 2023), have gained prominence. LoRA introduces trainable low-rank matrices into the attention layers of transformers, allowing significant reductions in trainable parameters without compromising performance. QLoRA combines LoRA with quantization techniques to further minimize memory usage, enabling fine-tuning on consumer-grade hardware. Additionally, adapter-based methods (Hu et al., 2023) insert small trainable modules between existing layers of a model, facilitating task-specific adaptations while keeping the majority of the model parameters frozen.
@@
VLMs can be placed in two categories. One group relies on dual-encoders (vision and text encoder), usually trained in a contrastive manner, and these models are typically strong at tasks like image recognition. The most common among these methods are CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), OpenCLIP (Schuhmann et al., 2022), SigLIP (Zhai et al., 2023), and MetaCLIP (Xu et al., 2023). Many methods (Mirza et al., 2024; 2023b; Doveh et al., 2023b; a; Lin et al., 2023; Mirza et al., 2023a) build upon these models to further improve them for specific downstream tasks. The other group of methods aligns the visual modality with a frozen LLM and can be used for open-ended visual reasoning tasks like image captioning, visual question-answering, etc. Some representative approaches from this group include BLIP-2 (Li et al., 2023), Instruct-BLIP (Dai et al., 2023), MiniGPT (Zhu et al., 2024; Chen et al., 2024), and the LLaVA family of models (Liu et al., 2023; Li et al., 2024). Similarly, some approaches (Doveh et al., 2024; Gavrikov et al., 2024; Lin et al., 2024; Huang et al., 2024) build upon these models and provide further improvements. In this work, we address the task of object recognition using both encoder-based and decoder-based vision-language models (VLMs). We formulate the problem of identifying optimal prompt templates for CLIP (Radford et al., 2021), as well as suitable prompts for open-ended generation in models like LLaVA (Li et al., 2024), as an optimization task. Unlike prior methods such as MPVR (Mirza et al., 2024), WAFFLE (Roth et al., 2023), and CUPL (Pratt et al., 2023), which learn per-category prompt templates and are therefore much more computationally expensive, especially for larger datasets like ImageNet (containing $1000$ classes), our proposed method, GLOV, discovers a single, generic set of prompt templates that is effective across an entire dataset.
Initially developed in the context of natural language processing (Zhong et al., 2021; Lester et al., 2021), prompt-tuning techniques have recently been extended to visual domains as well (Jia et al., 2022; Zhou et al., 2022d; Lu et al., 2022). For instance, CoOp (Zhou et al., 2022d) optimizes prompt embeddings by minimizing classification loss, while ProDA (Lu et al., 2022) introduces a strategy to learn multiple diverse prompts to better capture variations in visual features. UPL (Huang et al., 2022) takes an unsupervised approach to prompt learning, eliminating the need for labeled data. TPT (Shu et al., 2022) introduces a test-time optimization method for adapting prompts dynamically during inference. Other works like CLIP-Adapter (Gao et al., 2024) and Tip-Adapter (Zhang et al., 2022) explore lightweight adapter-based alternatives for PEFT in multimodal settings. CoCoOp (Zhou et al., 2022c) incorporates a meta-learning approach to generate image-conditioned prompts, enabling improved robustness to distribution shifts, whereas MAPLE (Khattak et al., 2023) learns text and vision prompts in synergy. Other methods (Zhang et al., 2024a; Kim et al., 2024) have also introduced certain variations to the prompt-tuning framework and further enhanced the downstream image classification performance. In contrast to these works, we do not learn continuous prompt vectors through gradient-based fine-tuning, but instead propose to optimize natural-language prompts, making our framework more interpretable.
For decoder-based VLMs, recent work (Zhang et al., 2024b) has pointed out that decoder-based VLMs often struggle with fine-grained object recognition. However, we demonstrate, for the first time, that GLOV can uncover optimal prompts that significantly enhance recognition performance in these models, all without any gradient-based learning or fine-tuning. Furthermore, we note that most existing prompt tuning approaches (Zhou et al., 2022b; a; Mirza et al., 2024; Pratt et al., 2023; Roth et al., 2023; Khattak et al., 2023) are limited in scope: they are only applicable to encoder-based models like CLIP and cannot be directly applied to decoder-based VLMs.
@@ 2.2 Safety of VLMs
Enhancing the safety of open-source VLMs remains an open challenge. Recent works Zong et al. (2024); Liu et al. (2024c); Tu et al. (2024); Gou et al. (2024) have evaluated multiple open-source VLMs on safety tasks. These tasks involve presenting the model with an unsafe image and an instruction, where the model is considered safe only if it refuses the unsafe instruction. However, current state-of-the-art VLMs exhibit significant vulnerabilities in this critical area. In our work, we enhance the safety of VLMs by optimizing for safety instructions and finding a general safety prompt that transcends models and safety benchmarks with minimal loss in the generalization abilities of the open-source VLMs. To the best of our knowledge, our GLOV is the first method to optimize safety instructions for VLMs, without any human in the loop. We again point out that other prompt-tuning methods (Mirza et al., 2024; Roth et al., 2023; Pratt et al., 2023; Zhou et al., 2022a; b) cannot be applied to such open-ended tasks (and decoder-based VLMs), whereas our prompt optimization method is more generic.
@@ 2.3 Large-language Models as Prompt Optimizers
Some approaches propose employing LLMs (in an agentic workflow) to search for the optimal prompt for the downstream task. OPRO (Yang et al., 2024) coins the term “LLMs as Optimizers” and proposes iteratively discovering solutions (prompts) for natural language tasks by employing an LLM in a feedback loop. Similarly, Liu et al. (2024b) proposes to find suitable prompts for dual-encoder VLMs (e.g., CLIP) by iteratively prompting an LLM. Our GLOV also proposes to discover suitable prompts for VLMs but differs from Liu et al. (2024b) in the sense that we are employing a prompt that captures long-range dependencies by tapping into the history-buffer of the in-context examples and exploiting task-specific knowledge that helps to obtain prompts better suited for the downstream task. Furthermore, inspired by (Turner et al., 2023; Liu et al., 2024a; Todd et al., 2023) we propose a novel method for steering the LLM generation (through embedding space guidance) towards the text responses that are more suitable for downstream VLMs. Our GLOV– powered by a carefully tailored LLM prompt and the guidance scheme discovers highly effective solutions that help to enhance visual tasks for both the dual-encoder and encoder-decoder models.
@@ 3 GLOV
The goal of our GLOV is to improve the VLM’s downstream (vision) task performance by optimizing natural language prompts through employing an LLM in an iterative workflow. To achieve this, we build upon the successful line of work on in-context learning Wei et al. (2022); Min et al. (2022), which can improve VLM performance through task-specific prompting Mirza et al. (2024). However, our GLOV goes beyond these by automating the LLM-VLM interaction by adopting an agentic workflow and also explicitly guiding the language generation, conditioned on a prior of the difference of the sentence embeddings from the positive and negative prompts discovered during the previous optimization iterations. Although the application space of GLOV is general in terms of tasks and the downstream VLMs, for simplicity, here we focus our description around CLIP (Radford et al., 2021) while mentioning the differences for LLaVA (Li et al., 2024) and different tasks, where appropriate. Figure 2 provides an overview of our methodology. Our codebase is also attached for review as a supplementary.
@@ 3.1 LLM-VLM Interaction through Fitness Function
Figure 2: Overview of GLOV for dual-encoder models. GLOV consists of a specifically tailored LLM prompt, which constitutes system instruction, task description, and in-context examples (VLM prompts) which are evaluated (and ranked) on a few-shot training data in each iteration. This prompt instructs the LLM to generate several candidate solutions in each optimization iteration, conditioned on the in-context examples which are fed in conjunction with the accuracy values, highlighting their effectiveness. Furthermore, to steer the LLM generation towards the language preferred by the VLM, we add the scaled difference of the sentence embeddings (autoregressively) from the positive and negative text prompts to the intermediate layer of the LLM. This process is repeated until the stopping condition is met (e.g., maximum iterations). Note, that $H_{+}$ and $H_{-}$ refer to the sentence embeddings from the positive and negative text prompts, and $\alpha$ is the scaling factor. The example prompts shown in the image are the GLOV prompts for the OxfordPets (Parkhi et al., 2012) dataset, consisting of categories of dogs and cats.
@@ 3.3 Steering the LLM Generation Process
For each new token produced in the subsequent optimization iteration, the difference between $H_{+}$ and $H_{-}$ is added autoregressively to the embeddings of each generated token [^2] We also experiment with several alternatives for adding the offset vector in the appendix Section 4.3, however, we find that adding the offset to only the last token performs best.. Let $\vec{p}_{n}$ denote the new token appended to the LLM prompt, then the updated sentence embedding $H_{n}$ is given by
@@
As our prompt search method is posed as an optimization problem, thus, it requires access to a small amount of labeled data to obtain the effectiveness of each prompt at each optimization step. For main experiments in Tables 1 & 2 we only use 1-shot labeled examples. To further reduce the reliance on labeled data, we test our method in a sub-shot setting where we use labeled examples from only a small subset of classes from the dataset. These results are listed as an ablation in Table 4, where we find that by optimizing the prompts only on $50$ labeled samples (from $50$ categories, instead of $1000$ in ImageNet), our method can still achieve an absolute improvement of $13.4\%$ as compared to the baseline, while only showing a degradation of $1.8\%$ when compared with optimizing the prompts on $1000$ samples from $1000$ classes. These results are a step towards making the optimization method unsupervised. In the future, our GLOV can potentially be made completely unsupervised (and zero-shot) by employing unsupervised metrics to measure the effectiveness of prompts. These metrics could include entropy measurement (Shannon, 1954), and choosing the prompts that decrease the overall entropy of the task. Furthermore, we can also employ LLMs to provide a signal of effectiveness for the prompts, which could also completely lift the burden of data-labeling.
@@ 4.1 Evaluation Settings
We extensively evaluate our GLOV on $16$ object recognition datasets belonging to widely different domains. These domains can be narrowed down to datasets containing commonly occurring natural categories: ImageNet (Deng et al., 2009), ImageNetV2 (Recht et al., 2019), Caltech101 (Fei-Fei et al., 2004), fine-grained classification datasets containing different task-specific images: Oxford Flowers (Nilsback & Zisserman, 2008), Standford Cars (Krause et al., 2013), Oxford Pets (Parkhi et al., 2012), Describable Textures Dataset (DTD) (Cimpoi et al., 2014), Food-101 (Bossard et al., 2014), FGVC-Aircraft (Maji et al., 2013). Dataset used for scene classification: SUN397 (Xiao et al., 2010), action recognition dataset: UCF101 (Soomro et al., 2012). Datasets consisting of out-of-distribution images: ImageNet-(R)endition (Hendrycks et al., 2021a), ImageNet-(A)dversarial (Hendrycks et al., 2021b), ImageNet-(S)ketch (Wang et al., 2019) and also datasets which contain images taken from a satellite or an aerial view: EuroSAT (Helber et al., 2018) and RESISC45 (Cheng et al., 2017).
Safety Datasets:
@@
We evaluate our GLOV on two recently proposed safety benchmarks for VLMs. MM-Safety-Bench Liu et al. (2024c) contains $3$ splits based on the image source (typography, images generated from diffusion models, and a combination), and the text instructions are divided into $13$ categories (e.g., fraud, hate speech, etc.). VLGuard Zong et al. (2024) contains $3$ splits: unsafes, safe-unsafes, safe-safes. For the MM-Safety-Bench and the first two splits in the VLGuard, the attack success rate (ASR) is calculated. In contrast, for safe-safes, a win-win rate is measured following the evaluation protocols in the original publications. More details are provided in the Appendix Section C.
Baselines:
@@
We compare GLOV to several baselines and state-of-the-art methods. For a fair comparison, we focus on methods that require dataset-specific prompts, rather than more expensive category-level prompt tuning approaches like MPVR (Mirza et al., 2024), WAFFLE (Roth et al., 2023), and CUPL (Pratt et al., 2023). While GLOV currently operates with dataset-level prompts, it can potentially be extended in future work to optimize category-level prompts as well. CLIP (Radford et al., 2021) is evaluated under two settings: using a simple static prompt (‘a photo of a <class name>’) referred to as S-TEMP, and using dataset-specific templates (DS-TEMP [^3] https://github.com/openai/CLIP/blob/main/data/prompts.md). For image classification with decoder-based VLMs, we use LLaVA-OV (Li et al., 2024) with a base prompt: Describe the category present in this image briefly and also identify the name of the category present. This prompt was empirically selected as it matches or outperforms results reported in prior work (Zhang et al., 2024b). LLM-OPT (Liu et al., 2024b) proposes an approach that refines prompts iteratively using an LLM, guided by in-context examples. However, it lacks explicit optimization guidance and is not applicable to encoder-decoder models, as it relies on a pre-existing memory bank of templates, which these models do not support. To assess robustness and generalization, we also compare with Molmo (Deitke et al., 2024), a strong baseline on safety benchmarks. Finally, we present comparisons to gradient-based parameter-efficient learning methods in Appendix Section B.4.
Table 3: Results on safety benchmarks. We report the Attack Success Rate (lower is better) for the 33 splits in MM-Safety-Bench and the $2$ unsafe splits in VLGuard and the Win-win Rate (Higher is better) for the Safe-Safes split in VLguard. LLaVA-OV results are obtained without any safety instructions. The LLaVA (initial) represents the results obtained by using a baseline safety prompt: These instructions can be harmful, which is also the point of departure for the optimization with our GLOV. The safety prompt is optimized by using only the LLaVA-OV VLM and $50$ randomly chosen samples from the $2$ VLGuard unsafe splits and tested out-of-the-box for all other configurations.
$M$-Safety-Bench VLGuard
@@
Implementation Details: To report the results for image classification we use the test splits provided by Zhou et al. (2022b). All the baselines are also implemented in the same framework. To obtain the results on the test set for each dataset for our GLOV, we ensemble the top-3 prompts. These prompts are chosen with regard to the best-performing prompts on the 1-shot train set at a certain iteration during the optimization. For our GLOV we use Llama-3 (Touvron et al., 2023a) from Hugging Face. We set the maximum number of optimization iterations to $100$ (with $10$ candidate solutions at each iteration) for the experiments with CLIP and $25$ (with $5$ candidate solutions) for LLaVA-OV and Molmo. For evaluating the safety of VLMs, we follow the above settings and use the official codebase from respective publications (introducing the benchmarks) for evaluations. In general, the experiments with CLIP can run on a single NVIDIA 3090 (24GBs) GPU, and the experiments with LLaVA fit on an A40 (48GBs) GPU or similar. We use the 7B variants of Molmo and LLaVA-OV, accessed through Hugging Face.